OpenStudy (anonymous):

sin2x = (2tanx)/(1+tan^2 x) only work on one side

OpenStudy (zarkon):

\[\frac{2\tan(x)}{1+\tan^2( x)}=\frac{2\tan(x)}{\sec^2(x)}=2\tan(x)\cos^2(x)=2\frac{sin(x)}{\cos(x)}\cos^2(x)\] \[=2\sin(x)\cos(x)=\sin(2x)\]